Ray Haskell Ford Lincoln Mercury

Inventory and Sales Report 2008 ESCAPE

History observed week of 12/31/2007 thru week of 04/14/2008 Dealership’s turn rate was 0% than surrounding market area.

The system will use the dealership sales mix rates to set inventory targets.

February 11, 2009

This page intentionally left blank 0 allocation

+Dealer 11607- 2008ESCAPE  1 of ??  February 11, 2009